OpenStudy (anonymous):

by logic, 4!= x3x2x1, 3!=3x2x1, why 1!=1 and 0!=1

OpenStudy (anonymous):

I think '0!=1' is by convention.

OpenStudy (anonymous):

yeh it's just one of the special rules

OpenStudy (anonymous):

a logical explanation please, why mathematicians decided this way

OpenStudy (anonymous):

they are humans, no gods, so I think there's a reason behind this

OpenStudy (anonymous):

would you accept the explanation n!= n*(n-1)! so if n=1 1!= 1*0! therefore 1=0!

OpenStudy (anonymous):

what about 0!=1

OpenStudy (anonymous):

its only where n is greater than or equal to 1 0!=1 is more like an assumption and there's really no actual way to prove it

OpenStudy (anonymous):

\[n!=n*(n-1)!\\\Rightarrow (n-1)!=\frac{n!}{n}\\\\\text{Substituting n=1,}\\\\(1-1)!=\frac{1!}{1}\\\\\Rightarrow 0!=1\]

OpenStudy (anonymous):

what about (-2)!

OpenStudy (anonymous):

there is no (-2)!

OpenStudy (anonymous):

if there is 0!, there is no reason theres no (-2)!

OpenStudy (anonymous):

just read the first three lines of: http://en.wikipedia.org/wiki/Factorial

OpenStudy (anonymous):

but theres no line saying negative numbers or non-integer(like 3.514) numbers CAN NOT have factorial..

OpenStudy (anonymous):

the very first line says "In mathematics, the factorial of a non-negative integer n, ......"

OpenStudy (anonymous):

what about the factorial of a negative number, or a decimal number.. what would they be

OpenStudy (anonymous):

for those who did some programming, I made a program using while loops that gives me the factorial of whatever number...

OpenStudy (anonymous):

and astonishingly..

OpenStudy (anonymous):

(please reply if u have any programming experience with loops...)

OpenStudy (anonymous):

astonishingly would give you a complex infinity error if your input for n is a negative number

OpenStudy (anonymous):

nop..

OpenStudy (anonymous):

far from that...

OpenStudy (anonymous):

when i did factorial of 1, it gave me 1... and when i did factorial of 0, it actually gave me 1 too..

OpenStudy (anonymous):

n! is a sequence with integer values for nonnegative n

OpenStudy (anonymous):

and factorial of any negative number gives also 1...

OpenStudy (anonymous):

well the rule states nonnegative integer. so ZERO is included.

OpenStudy (anonymous):

did you create a function loop that when the input for n is a negative value that the answer would yield 1?

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

the program is written in python, through pure logic, no exceptions are defined

OpenStudy (anonymous):

written in python: def factorial(n): i=1 a=1 while i<=n: a=a*i i=i+1 return a

OpenStudy (anonymous):

your while i<=n statement returns a value of 1

ganeshie8 (ganeshie8):

try this : def factorial(n): i=1 a=1 while i<=n: a=a*i i=i+1 return a print "BAD INPUT !"

OpenStudy (anonymous):

only when n is less than 1.9999999

OpenStudy (anonymous):

@ganeshie8 your return a has bad incantation, it should be placed after the loop is completely finished

OpenStudy (anonymous):

\[n!\neq0/;\forall n\]

ganeshie8 (ganeshie8):

yeah my code will not work as well :( i think your code is not entering while loop at all for negative integers. it is simply returning the value of a

OpenStudy (anonymous):

apply that zero rule

OpenStudy (anonymous):

yes ganesh. it's just going to print or display 1 when a negative value is the input

OpenStudy (anonymous):

factorial in programming has more than 100 formulas

OpenStudy (anonymous):

\[! == Product[k, {k, 1, n}] /; Element[n, \\mathbb{Z}s] && n >= 1\]

OpenStudy (anonymous):

but what's truly fascinating is.. when you program through pure logic, as you can see in the program, it logically multiplies every numbers from the input, descending til 1. And when you do that, if you enter 1 as the input, it gives 1. If u enter 0 as input, normally through the logic, it would be absurd, but the program still gives 1

OpenStudy (anonymous):

which perfectly fits the mathematical rules that are built centuries ago, when machines didn't exist yet

OpenStudy (anonymous):

but zero is 1 secondary to empty product rule

OpenStudy (anonymous):

also, have you looked at the list prod function in python?

OpenStudy (anonymous):

'but zero is 1 secondary to empty product rule' what do u mean by that

OpenStudy (anonymous):

what do you get when you raise any number (n) to zero? with the exception of n=0

OpenStudy (anonymous):

1

OpenStudy (anonymous):

then what is zero raised to zero?

OpenStudy (anonymous):

if you're programming, you can set 0^0=1 when representing an empty product

OpenStudy (anonymous):

if u program an exponential function using a pure logical algorithm, i think there is no need to add an exception, everything will just display as it actually is mathematically

OpenStudy (anonymous):

on what logic would you set the domain and range of the function then?

OpenStudy (anonymous):

there's no need to set them, with any input, the output will never reach 0 or below

OpenStudy (maheshmeghwal9):

There is a negative factorial but not for negative integers. The factorial function can be extended so that it is defined for all real numbers (and complex numbers) except for negative integers. It is extended by using the gamma function. (see source). Source(s): http://en.wikipedia.org/wiki/Factorial

OpenStudy (anonymous):

how would you tackle the problem of providing the factorial of multiple elements in a set?

OpenStudy (anonymous):

ie: {2, 3, 5, 7, 11}!

OpenStudy (anonymous):

what does this mean? I only know what factorial is

OpenStudy (anonymous):

the factorial of {2, 3, 5, 7, 11}

OpenStudy (anonymous):

or series expansion at infinity

OpenStudy (anonymous):

give me an example please, I'm not sure what you mean by a factorial of multiple elements in a set

OpenStudy (anonymous):

I just gave you one

OpenStudy (anonymous):

{2, 3, 5, 7, 11}!

OpenStudy (anonymous):

yes, what does {2, 3}! mean?

OpenStudy (anonymous):

2x1x3x2x1 or what..

OpenStudy (anonymous):

btw, I just made a simple python program using PURE LOGIC, no exceptions are added, and computer seems to behave the same way as the actual mathematics, hence, 0^0 gives one def power(x,n): a = 1 i = 0 while i < n: a = a * x i = i + 1 return a

OpenStudy (anonymous):

nice discussion with you panlac01, have to go sleep mate, we can talk about it tomorrow ;)

Parth (parthkohli):

I know this one! \((n - 1)! = \Large {n! \over n}\)